# Orbital decay candidates reconsidered: WASP-4 b is not decaying and Kepler-1658 b is not a planet

### v3

@@ Abstract
The fate of hot Jupiters is thought to be engulfment by their host stars, the outcome of tidal orbital decay. Transit timing has revealed a few systems with apparently shrinking orbital periods, but such signals can be mimicked by light travel-time effects (LTTE) of a distant companion. Combining transit timings with precise radial-velocity data, including new data, we reassessed three reported cases of orbital decay: WASP-4, WASP-12, and Kepler-1658. For WASP-4, the period change is best explained by LTTE due to an $\approx$$8~M_{\rm Jup}$ companion at $\approx$8 AU, with no need to invoke orbital decay. For WASP-12, in contrast, the data firmly exclude LTTE and confirm genuine orbital decay. For Kepler-1658, spectroscopic and photometric anomalies reveal the “planet” to be an eclipsing K/M binary bound to the F-type primary, with LTTE explaining the observed period change. Thus, among the known hot Jupiters, only WASP-12 b currently shows compelling evidence for orbital decay.
@@ 1 Introduction
Hot Jupiters have famously small orbits, some of which are only 3-4 times larger than their host stars. These hottest of the hot Jupiters are expected to be relatively short-lived on cosmological timescales, slowly spiraling into their host stars as tidal interactions sap their orbital energy and angular momentum. In close stellar binaries, tidal interactions usually lead to spin-orbit synchrony (Hut, 1980), but a hot Jupiter’s angular momentum is insufficient to spin a Sun-like star into synchrony (Rasio et al., 1996; Levrard et al., 2009), especially since magnetized winds steadily slow the star’s rotation (Barker & Ogilvie, 2009). However, the timescale of tidal orbital decay remains uncertain because it depends on poorly understood mechanisms of tidal dissipation within stars, such as inertial-wave damping and turbulent convection (see, e.g., Ogilvie, 2014).
@@ 2 Mathematical preliminaries
where $E$ is the eccentric anomaly, $t_{\rm p}$ is the time of periapse passage, and $n$ is the mean motion ($2\pi$ divided by the orbital period $P$). By taking the time derivative, one obtains the radial-velocity equation,
@@
A few changes in parameterization are useful. First, we rewrite $a\sin I$ in terms of the radial-velocity semi-amplitude,
@@
because $K$ is more directly observable. Second, instead of using $t_{\rm p}$ to pinpoint the star’s location at a particular time, we use the true anomaly at inferior conjunction,
@@
where $v_{0}$ is the radial velocity of the planetary system’s center of mass relative to the Solar System, $\gamma$ is a spectrograph-dependent additive offset [^1] The $\gamma$ parameter is necessary because spectrographs measure radial-velocity variations at the m/s level, but the accuracy of the absolute radial velocity scale is usually no better than 100 m/s, leading to systematic offsets between different spectrographs., and $\dot{z}_{1}$ and $\dot{z}_{2}$ are the contributions from the hot Jupiter and wider-orbiting companion, respectively.
@@
$t_{k}\approx t_{0}+k\left(\!P+\frac{\beta_{0}}{c}\!\right)+\frac{1}{c}\sum_{j=1}^{\infty}\beta_{j}\frac{(t_{0}+kP)^{j+1}-t_{0}^{j+1}}{j+1},$ (14)
@@ 3 WASP-4
Whether or not orbital decay has been detected for WASP-4 b has been the subject of debate. Early timing studies showed the transit period to be shrinking at a rate of order $\delta P/P=10~{\rm ms/yr}$ (Bouma et al., 2019; Southworth et al., 2019), although the robustness of this finding was questioned (Baluev et al., 2019). Additional RV data suggested the star is accelerating a rate compatible with the LTTE hypothesis (Bouma et al., 2020), but concerns were raised about systematic errors (Baluev et al., 2020). With additional data, Turner et al. (2022) confirmed a shrinking period and argued that the LTTE hypothesis was firmly ruled out, but an error in their analysis was later identified, showing that LTTE was still viable (Harre & Smith, 2023). Most recently, Baştürk et al. (2025) reported further transit times and concluded that orbital decay is the best explanation.
@@ 3.1 Radial Velocities
  • We added a new HIRES velocity from December 26, 2020 (BJD 2459189.71567), extending the HIRES time span by 1.3 years.
  • To mitigate the effects of correlated errors due to stellar rotation (Husnoo et al., 2012), we “thinned” the data by retaining only one point every five days.
@@ 3.2 Transit Times
WASP-4 was observed in Sectors 2, 28, 29, 69, and 96 of NASA’s Transiting Exoplanet Survey Satellite (TESS) mission (Ricker et al., 2015), and all the data are available with 120 s time sampling. The Sector 96 data have not previously been analyzed in the literature. We used light curves processed by the TESS Science Processing and Operations Center (SPOC) to measure transit times with the procedures of Ivshina & Winn (2022), obtaining 70 transit times with typical uncertainties of 25 s.
@@
  • include the pre-TESS transit times compiled by Ivshina & Winn (2022), which are drawn exclusively from the peer-reviewed literature,
@@
  • discard the three data points with formal uncertainties $\geq$60 s, and
@@
Figure 1: WASP-4. Upper left.—Radial velocities. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=1.338231293$ days, $t_{\rm c}=2456180.558583$ BJDTDB). In this panel and others, the black curve is the best-fit model including the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time), after subtracting the best-fit contribution from the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
@@ 3.3 Timing+RV analysis
The model had nine free parameters: $K_{1}$, $P_{1}$, $t_{\rm c,1}$, the three $\beta$ parameters, and the three $\gamma$ offsets (one per spectrograph). The best-fit yielded $\chi^{2}=184$ with 164 degrees of freedom, indicating an acceptable fit. Posterior sampling was performed using the Markov Chain Monte Carlo (MCMC) algorithm of Goodman & Weare (2010) as implemented in emcee (Foreman-Mackey et al., 2013). To allow for possible underestimation of uncertainties, we introduced a velocity jitter term for each spectrograph. Table 1 gives the results and Figure 1 shows the corresponding fits. The model provides a good fit to all the observables.
@@
$K_{1}$ $237.6$ $237.6^{+2.7}_{-2.7}$ m/s
$\beta_{1}$ 17.1-17.1 $17.1^{+1.2}_{-1.3}$ m/s/yr
$\beta_{2}$ 1.05-1.05 $1.06^{+0.17}_{-0.16}$ m/s/yr2
$\beta_{3}$ $0.171$ $0.174^{+0.041}_{-0.041}$ m/s/yr3
$\gamma_{\rm CORALIE}$ 35.2-35.2 $35.1^{+6.8}_{-6.8}$ m/s
$\gamma_{\rm HARPS}$ 63.8-63.8 $63.3^{+8.5}_{-9.2}$ m/s
$\gamma_{\rm HIRES}$ 46.2-46.2 $46.1^{+3.5}_{-3.7}$ m/s
$\sigma_{\rm CORALIE}$ $10.1$ $10.1^{+6.7}_{-6.3}$ m/s
$\sigma_{\rm HARPS}$ $11.5$ $11.5^{+9.6}_{-4.9}$ m/s
$\sigma_{\rm HIRES}$ $5.3$ $5.3^{+2.9}_{-2.0}$ m/s
$a_{2}$ $8.39$ $7.95^{+0.98}_{-0.41}$ AU
$P_{2}$ $25.6$ $23.6^{+4.5}_{-1.9}$ yr
$K_{2}$ $78.1$ $73.6^{+12.2}_{-4.7}$ m/s
$m_{2}\sin I_{2}$ $7.56$ $6.94^{+1.59}_{-0.60}$ $M_{\rm Jup}$
@@ 4 WASP-12
Figure 3: WASP-12. Upper left.—Radial velocities. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=1.091418667$ days, $t_{\rm c}=2456305.45552$ BJDTDB). In this panel and others, the black curve is the best-fit model including the effects of the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time), after subtracting the best-fit contribution of the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
@@
WASP-12 has been classified as 6500 K main-sequence star with mass 1.4 $M_{\odot}$ and 1.7 $R_{\odot}$ (Collins et al., 2017; Bailey & Goodman, 2019), although there are suggestions it may be somewhat less massive and more evolved (Weinberg et al., 2017). The hot Jupiter WASP-12 b, discovered by Hebb et al. (2009), has mass 1.5 $M_{\rm Jup}$, radius $1.9~R_{\rm Jup}$, and period 1.1 days. With an orbital separation only three times the stellar radius, WASP-12 b lies precariously close to the Roche limit (Antonetti & Goodman, 2022). We do not attempt to review the vast WASP-12 literature (see Haswell, 2018, for an overview). Our aim was to apply the same joint timing + RV analysis used for WASP-4 to test the viability of the LTTE hypothesis for this system.
@@
TESS data with 120 s sampling are available from Sectors 20, 43, 44, 45, 71, and 72. We re-analyzed the SPOC light curves and combined the measured transit times with the pre-TESS compilation of Ivshina & Winn (2022). As before, we excluded ground-based data obtained after the first TESS observation, imposed a minimum uncertainty of 25 s and discarded times with formal uncertainties exceeding 60 s.
@@
There are three prior sources of radial velocities: the SOPHIE [^5] SOPHIE is the Spectrographe pour l’Observation des Phénomènes des Intérieurs stellaires et des Exoplanètes, mounted on the 1.9 m telescope at Haute-Provence Observatory, France. data of Hebb et al. (2009), the HARPS-N [^6] HARPS-N is the High Accuracy Radial velocity Planet Searcher for the Northern hemisphere, fed by the 3.6 m Telescopio Nazionale Galileo at Roque de los Muchachos Observatory on the La Palma, Canary Islands, Spain. data of Bonomo et al. (2017), and the HIRES data of Yee et al. (2020). We supplemented these with 20 new velocities obtained with the NEID [^7] NEID stands for NN-EXPLORE Exoplanet Investigations with Doppler Spectroscopy. It is a stabilized optical spectrograph with resolution $\approx$120,000 (Schwab et al., 2016) mounted on the WIYN 3.5 m telescope at Kitt Peak, Arizona. spectrograph in 2024 and 2025. The NEID spectra were processed with version 1.4.2 of the data reduction pipeline. [^8] https://neid.ipac.caltech.edu/docs/NEID-DRP/ Spectra obtained during transits were excluded, and the time series was thinned to enforce a five-day minimum time spacing.
@@
To test the LTTE hypothesis, we used a model similar to that used for WASP-4, modified only to account for the RV anomalies caused by the star’s tidal distortion (Arras et al., 2012). For small distortions, the effect introduces a sinusoidal signal with a period of $P_{1}/2$, an amplitude of a few m/s, and a phase shift of 90 relative to the main signal. We modeled this effect by the simple expedient of assigning the planet a fictitious eccentricity with $\omega=-\pi/2$ (for more details, see Arras et al., 2012; Maciejewski et al., 2020).
The best-fit model is shown in Figure 3 (dashed black curves). The fit is very poor, with with $\chi^{2}=1{,}531$ and 323 degrees of freedom. The model attempts to reproduce the quadratic timing trend with a suitably strong line-of-sight acceleration (lower left) but is stymied by the constraint imposed by contemporaneous RV data (lower right), especially the HIRES data (green points).
@@ 5 Kepler-1658
Figure 4: Kepler-1658. Upper left.—Radial velocities. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=3.849363145$ days, $t_{\rm c}=2455005.925555$ BJDTDB). In this panel and others, the black curve is the best-fit model including the effects of the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time) after subtracting the best-fit contribution of the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
@@
Vissapragada et al. (2022) analyzed TESS data from Sectors 41, 54, and 55, as well as two ground-based observations with the Wide Field Infrared Camera (WIRC) on the Hale 5 m telescope at Palomar Observatory, California. When combined with the Kepler data, these transit times revealed apparent orbital decay at a rate of 130 ms/yr, five times faster than WASP-12 b and greatly exceeding theoretical expectations (Barker et al., 2024).

### v4

@@ Abstract
The fate of hot Jupiters is thought to be engulfment by their host stars, the outcome of tidal orbital decay. Transit timing has revealed a few systems with apparently shrinking orbital periods, but such signals can be mimicked by light travel-time effects (LTTE) of a distant companion. By combining transit timings with precise radial-velocity data, including new data, we reassessed three reported cases of orbital decay: WASP-4, WASP-12, and Kepler-1658. For WASP-4, the period change is best explained by LTTE due to a $\approx$$7~M_{\rm Jup}$ companion at $\approx$8 AU, with no need to invoke orbital decay. For WASP-12, in contrast, the data firmly exclude LTTE and confirm genuine orbital decay. For Kepler-1658, spectroscopic and photometric anomalies reveal the “planet” to be an eclipsing K/M binary bound to the F-type primary, with LTTE explaining the observed period change. Thus, among the known hot Jupiters, only WASP-12 b currently shows compelling evidence for orbital decay.
@@ 1 Introduction
Hot Jupiters have famously small orbits, some of which are only 3-4 times larger than their host stars. These hottest of the hot Jupiters are expected to be relatively short-lived on cosmological timescales, slowly spiraling into their host stars as tidal interactions sap their orbital energy and angular momentum. In close stellar binaries, tidal interactions usually lead to spin-orbit synchronization (Hut, 1980), but a hot Jupiter’s angular momentum is insufficient to spin a Sun-like star into synchrony (Rasio et al., 1996; Levrard et al., 2009), especially since magnetized winds steadily slow the star’s rotation (Barker & Ogilvie, 2009). However, the timescale of tidal orbital decay remains uncertain because it depends on poorly understood mechanisms of tidal dissipation within stars, such as inertial-wave damping and turbulent convection (see, e.g., Ogilvie, 2014).
@@ 2 Mathematical preliminaries
where $E$ is the eccentric anomaly, $t_{\rm p}$ is the time of periapse passage, and $n$ is the mean motion ($2\pi$ divided by the orbital period $P$). By taking the time derivative of Equation 1, one obtains the radial-velocity equation,
@@
A few changes in parameterization are useful. First, we define the radial-velocity semi-amplitude,
@@
because it can be measured directly from radial-velocity data. Second, instead of using $t_{\rm p}$ to pinpoint the star’s location at a particular time, we use the true anomaly at inferior conjunction,
@@
where $v_{0}$ is the radial velocity of the planetary system’s center of mass relative to the Solar System, $\gamma$ is a spectrograph-dependent additive offset [^1] The $\gamma$ parameter is necessary because although spectrographs can measure radial-velocity variations with m/s precision, the accuracy of the absolute radial velocity scale is usually no better than 100 m/s, leading to systematic offsets between different spectrographs., and $\dot{z}_{1}$ and $\dot{z}_{2}$ are the contributions from the hot Jupiter and wider-orbiting companion, respectively.
@@
$t_{k}\approx t_{0}+k\left(\!P_{1}+\frac{\beta_{0}}{c}\!\right)+\frac{1}{c}\sum_{j=1}^{\infty}\beta_{j}\frac{(t_{0}+kP_{1})^{j+1}-t_{0}^{j+1}}{j+1},$ (14)
@@ 3 WASP-4
Whether or not orbital decay has been detected for WASP-4 b has been the subject of debate. Early timing studies showed the transit period to be shrinking at a rate of order $\delta P/P=10~{\rm ms/yr}$ (Bouma et al., 2019; Southworth et al., 2019), although the robustness of this finding was questioned (Baluev et al., 2019). Additional RV data suggested the star is accelerating at a rate compatible with the LTTE hypothesis (Bouma et al., 2020), but concerns were raised about systematic errors (Baluev et al., 2020). With additional data, Turner et al. (2022) confirmed a shrinking period and argued that the LTTE hypothesis was firmly ruled out, but an error in their analysis was later identified, showing that LTTE was still viable (Harre & Smith, 2023). Most recently, Baştürk et al. (2025) reported further transit times and concluded that orbital decay is the best explanation.
@@ 3.1 Radial Velocities
  • We added a new HIRES velocity from December 26, 2020 (BJD 2459189.71567), extending the HIRES time span by 1.3 years. The full HIRES dataset is given in the Appendix (Table 2).
  • We “thinned” the data by retaining only one point every five days. This was done to avoid the complexity of modeling time-correlated noise. [^5] Husnoo et al. (2012) fitted RV data with a model in which the correlation strength is a Gaussian function of the time interval between observations, and for WASP-4 the Gaussian width was 1.5 days. Thus, points separated 5 days are expected to have minimal correlations.
@@ 3.2 Transit Times
WASP-4 was observed in Sectors 2, 28, 29, 69, and 96 of NASA’s Transiting Exoplanet Survey Satellite (TESS) mission (Ricker et al., 2015), and all the data are available with 120 s time sampling. The Sector 96 data have not previously been analyzed in the literature. We used light curves processed by the TESS Science Processing and Operations Center (SPOC) to measure transit times with the procedures of Ivshina & Winn (2022a), obtaining 78 transit times with typical uncertainties of 25 s. The new transit times from Sector 96 are given in the Appendix (Table 3).
@@
  • include the pre-TESS transit times compiled by Ivshina & Winn (2022a), which are drawn exclusively from the peer-reviewed literature,
@@
  • discard the three data points with formal uncertainties exceeding 60 s, and
@@
Figure 1: WASP-4. Upper left.—Radial velocities, Radial velocities, after shifting the data from each spectrograph to have zero mean. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=1.338231293$ days, $t_{\rm c}=2456180.558583$ BJDTDB). In this panel and others, the black curve is the best-fit model including the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time), after subtracting the best-fit contribution from the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
@@ 3.3 Timing+RV analysis
The model had nine free parameters: $K_{1}$, $P_{1}$, $t_{\rm c,1}$, the three $\beta$ parameters, and the three $\gamma$ offsets (one per spectrograph). The best-fit model yielded $\chi^{2}=187.8$ with 172 degrees of freedom, indicating an acceptable fit. Posterior sampling was performed using the Markov Chain Monte Carlo (MCMC) algorithm of Goodman & Weare (2010) as implemented in emcee (Foreman-Mackey et al., 2013). To allow for possible underestimation of uncertainties, we introduced a velocity jitter term for each spectrograph. Table 1 gives the results and Figure 1 shows the corresponding fits. The model provides a good fit to all the observables.
@@
$K_{1}$ $237.6$ $237.6^{+2.6}_{-2.6}$ m/s
$\beta_{1}$ 17.4-17.4 $17.3^{+1.0}_{-1.1}$ m/s/yr
$\beta_{2}$ 1.03-1.03 $1.04^{+0.15}_{-0.14}$ m/s/yr2
$\beta_{3}$ $0.180$ $0.181^{+0.033}_{-0.032}$ m/s/yr3
$\gamma_{\rm CORALIE}$ 35.7-35.7 $35.2^{+6.8}_{-6.8}$ m/s
$\gamma_{\rm HARPS}$ 64.5-64.5 $63.6^{+8.5}_{-9.1}$ m/s
$\gamma_{\rm HIRES}$ 46.9-46.9 $46.6^{+3.1}_{-3.4}$ m/s
$\sigma_{\rm CORALIE}$ $10.1$ $10.2^{+6.7}_{-6.3}$ m/s
$\sigma_{\rm HARPS}$ $11.5$ $11.6^{+9.6}_{-4.9}$ m/s
$\sigma_{\rm HIRES}$ $5.3$ $5.2^{+2.9}_{-2.0}$ m/s
$a_{2}$ $8.29$ $7.96^{+0.72}_{-0.32}$ AU
$P_{2}$ $25.2$ $23.7^{+3.3}_{-1.4}$ yr
$K_{2}$ $76.9$ $73.4^{+8.6}_{-3.8}$ m/s
$m_{2}\sin I_{2}$ $7.39$ $6.93^{+1.10}_{-0.47}$ $M_{\rm Jup}$
@@ 4 WASP-12
Figure 3: WASP-12. Upper left.—Radial velocities, after shifting the data from each spectrograph to have zero mean. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=1.091418667$ days, $t_{\rm c}=2456305.45552$ BJDTDB). In this panel and others, the black curve is the best-fit model including the effects of the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time), after subtracting the best-fit contribution of the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
@@
WASP-12 has been classified as 6500 K main-sequence star with mass 1.4 $M_{\odot}$ and radius 1.7 $R_{\odot}$ (Collins et al., 2017; Bailey & Goodman, 2019), although there are suggestions it may be somewhat less massive and more evolved (Weinberg et al., 2017). The hot Jupiter WASP-12 b, discovered by Hebb et al. (2009), has mass 1.5 $M_{\rm Jup}$, radius $1.9~R_{\rm Jup}$, and period 1.1 days. With an orbital separation only three times the stellar radius, WASP-12 b lies precariously close to the Roche limit (Antonetti & Goodman, 2022). We do not attempt to review the vast WASP-12 literature (see Haswell, 2018, for an overview). Our aim was to apply the same joint timing + RV analysis used for WASP-4 to test the viability of the LTTE hypothesis for this system.
@@
TESS data with 120 s sampling are available from Sectors 20, 43, 44, 45, 71, and 72. We re-analyzed the SPOC light curves and combined the measured transit times with the pre-TESS compilation of Ivshina & Winn (2022a). As before, we excluded ground-based data obtained after the first TESS observation, imposed a minimum uncertainty of 25 s and discarded times with formal uncertainties exceeding 60 s.
@@
There are three prior sources of radial velocities: the SOPHIE [^6] SOPHIE is the Spectrographe pour l’Observation des Phénomènes des Intérieurs stellaires et des Exoplanètes, mounted on the 1.9 m telescope at Haute-Provence Observatory, France. data of Hebb et al. (2009), the HARPS-N [^7] HARPS-N is the High Accuracy Radial velocity Planet Searcher for the Northern hemisphere, fed by the 3.6 m Telescopio Nazionale Galileo at Roque de los Muchachos Observatory on the La Palma, Canary Islands, Spain. data of Bonomo et al. (2017) and Maciejewski et al. (2020), and the HIRES data of Yee et al. (2020). We supplemented these with 20 new velocities obtained with the NEID [^8] NEID stands for NN-EXPLORE Exoplanet Investigations with Doppler Spectroscopy. It is a stabilized optical spectrograph with resolution $\approx$120,000 (Schwab et al., 2016) mounted on the WIYN 3.5 m telescope at Kitt Peak, Arizona. spectrograph in 2024 and 2025. The NEID spectra were processed with version 1.4.2 of the data reduction pipeline. [^9] https://neid.ipac.caltech.edu/docs/NEID-DRP/ Spectra obtained during transits were excluded, and the time series was thinned to enforce a five-day minimum time spacing. The resulting velocities are given in the Appendix (Table 4). To account for stellar activity, we adopted a velocity jitter of 10 m/s, the approximate value determined in single-planet fits over time intervals less than a few years (Bonomo et al., 2017; Maciejewski et al., 2020; Yee et al., 2020).
@@
To test the LTTE hypothesis, we used a model similar to that used for WASP-4, modified only to account for the RV anomalies caused by the star’s tidal distortion (Arras et al., 2012). For small distortions, the effect introduces a sinusoidal signal with a period of $P_{1}/2$, an amplitude of a few m/s, and a phase shift of 90 relative to the main signal. We modeled this effect by the simple expedient of assigning the planet a fictitious eccentricity and fixing $\omega=-\pi/2$ (for more details, see Arras et al., 2012; Maciejewski et al., 2020).
The best-fit model is shown in Figure 3 (dashed black curves). The fit is very poor, with $\chi^{2}=1{,}531$ and 323 degrees of freedom. The model attempts to reproduce the quadratic timing trend with a suitably strong line-of-sight acceleration (lower left) but is stymied by the constraint imposed by contemporaneous RV data (lower right), especially the HIRES data (yellow diamonds).
@@ 5 Kepler-1658
Figure 4: Kepler-1658. Upper left.—Radial velocities, after shifting the data from each spectrograph to have zero mean. Lower left.—Deviations between transit times and the best-fit constant-period model ($P=3.849363145$ days, $t_{\rm c}=2455005.925555$ BJDTDB). In this panel and others, the black curve is the best-fit model including the effects of the hot Jupiter and a wide-orbiting companion, and the gray curves are random draws from the posterior. Upper right.—Radial velocities as a function of folded time (i.e. relative to the nearest transit time) after subtracting the best-fit contribution of the wide-orbiting companion. Lower right.—Radial velocities after subtracting the best-fit contribution of the hot Jupiter.
@@
Vissapragada et al. (2022) analyzed TESS data from Sectors 41, 54, and 55, as well as two ground-based observations with the Wide Field Infrared Camera (WIRC) on the Hale 5 m telescope at Palomar Observatory, California. When combined with the Kepler data, these transit times revealed apparent orbital decay at a rate of 130 ms/yr, five times faster than for WASP-12 b and greatly exceeding theoretical expectations (Barker et al., 2024).